### v1

@@ 3.2 Parameter inference
The multi-band observational data and the model predictions based on the best-fit parameters are shown in Figure 7. The FS+RS afterglow model provides a good explanation for the optical data and is roughly consistent with the GeV observations (despite large uncertainties), but it does not explain the X-ray observations well, suggesting that there may be an additional contribution to the X-ray emission. By combining the afterglow model and prompt emission to constrain the initial Lorentz factor, we can examine the tight parameter correlation proposed by Liang et al. (2015), i.e. . As illustrated in the left side of Figure 8, GRB 240825A generally aligns with this correlation and, similar to other GRBs with RS emission (hereafter referred to as RS GRBs) (Japelj et al. 2014), exhibits higher luminosity compared to other typical Type II GRBs. Furthermore, based on the microphysics parameters of magnetic energy in the forward shock and reverse shock obtained from the posterior parameters, we can estimate the magnetic ratio of RS and FS jet with 171717 is defined in Zhang & Kobayashi (2005). To compare with some statistical studies (Japelj et al. 2014; Yi et al. 2020; Wang et al. 2024), we adopt here, which requires taking the square root for conversion to the magnetic field ratio.. As shown on the right side of Figure 8, GRB 240825A exhibits a RS-to-FS magnetic ratio that is in the high end of the distribution derived by (Wang et al. 2024). And the of host galaxy we finally obtained as free parameters matches well the constraint obtained by Cheng et al. (2025).
@@ 4 Summary and Discussion
In this work, we performed a detailed analysis of the prompt emission spectrum using data from Swift-BAT and Fermi-GBM/LAT and modeled the multi-wavelength (optical/NIR, X-ray, -ray) afterglow light curves using the FS+RS model. By combining the observational data with our analysis results, we summarize the characteristics of this event as follows:
The prompt emission spectrum of GRB 240825A is best fit by a model combining a Band function and a power-law component, with a high-energy cutoff is 76 MeV. The peak energy () of the Band function is approximately 500 keV, and the -corrected isotropic energy is estimated to be 1.7 erg. These values make GRB 240825A consistent with the Amati relation.
Considering the high-energy cutoff of the prompt emission due to absorption and the radiation efficiency ( 0.1), constraints yield an initial Lorentz factor 245 and an isotropic kinetic energy 1.7 erg. These two constraints from prompt emission were used to update the prior ranges of parameters during Bayesian inference in the afterglow modelling.
@@
Based on the aforementioned characteristics of GRB 240825A, we have conducted discussions on the following points. Firstly, GRB 240825A is an event with comprehensive observational coverage across time and wavelength, spanning from the prompt emission to the afterglow phase. During the afterglow phase, the emission is initially dominated by RS components, and transitions to being FS dominated, and eventually reflects contributions from the host galaxy. Secondly, such a “flattening case” in the combination of FS+RS typically requires 1 (Zhang et al. 2003), which is consistent with the result obtained from our model fitting. This suggests that the RS is more magnetized than the FS, but is not Poynting-flux-dominated (otherwise the RS is suppressed). Secondly, although this burst appears to be a typical Type II GRB based on empirical correlations, its comprehensive observational data make it an excellent probe for studying the physics of GRB jets. For example, the initial Lorentz factor of the afterglow is often correlated with the onset time. Although SVOM/C-GFT conducted rapid follow-up observations of GRB 240825A, the optical rise phase is still missing. Therefore, the constraints on the Lorentz factor derived from the prompt emission are crucial, highlighting the importance of comprehensive observations from the prompt emission to the afterglow phase. Thirdly, despite the overall consistency between the model and the observed multi-wavelength afterglow data, the X-ray light curve is not well reproduced in our current framework. This discrepancy may point to additional physical components or emission mechanisms not accounted for in the standard external shock model. Several scenarios have been proposed in the literature to explain the shaping of X-ray afterglow emission in GRBs, such as energy injection, evolving shock parameters, etc. (Zhang et al. 2006; Fan & Piran 2006). However, most of these scenarios generally predict correlated features in other bands, particularly in the optical or high-energy regimes. The fact that our model provides a good fit to the optical, while significantly underpredicting the X-ray flux, suggests that the X-ray excess may arise from an additional component that selectively contributes to the X-ray band. In previous work, such as GRB 130427A, X-rays may come from the contributions of flash-RS and RS Vestrand et al. (2014). In addition, based on GRB 221009A, people realized that there may be an extremely narrow component in the jet structure Zhang et al. (2024); Ren et al. (2024), i.e., the two-component jet model. Hence, this remains an open question, and future observations and modeling efforts will be necessary to fully understand the origin of such X-ray behavior.
@@ Acknowledgements.
The Space-based multi-band Variable Objects Monitor (SVOM) is a joint Chinese-French mission led by the Chinese National Space Administration (CNSA), the French Space Agency (CNES), and the Chinese Academy of Sciences (CAS). We gratefully acknowledge the unwavering support of NSSC, IAMCAS, XIOPM, NAOC, IHEP, CNES, CEA, and CNRS. This work is supported by the National Key R&D Program of China (Grant No. 2024YFA1611600), the SVOM project (a mission under the Strategic Priority Program on Space Science of the Chinese Academy of Sciences), the Strategic Priority Research Program of the Chinese Academy of Sciences (Grant No. XDB0550401), and the National Natural Science Foundation of China (Grant No. 12494573). This research was partly supported by Natural Science Foundation of Xinjiang Uygur Autonomous Region (Grant No. 2024D01D32), Tianshan Talent Training Program (Grant No. 2023TSYCLJ0053), and Tianshan Innovation Team Program (Grant No. 2024D14015). This work is also supported by the following funding programs: Y.W. is supported by the Jiangsu Funding Program for Excellent Postdoctoral Talent (Grant No. 2024ZB110), the Postdoctoral Fellowship Program (Grant No. GZC20241916), and the General Fund (Grant No. 2024M763531) of the China Postdoctoral Science Foundation. G.P.L. is supported by a Royal Society Dorothy Hodgkin Fellowship (grant Nos. DHF-R1-221175 and DHF-ERE-221005). A.S. acknowledges support by a postdoctoral fellowship from the CNES. X.H.H. is supported by the National Key R&D Program of China (Grant No. 2024YFA1611702) and the Strategic Priority Research Program of the Chinese Academy of Sciences (Grant No. XDB0550101). S.D.V., E.L.F., B.S. acknowledge the support of the French Agence Nationale de la Recherche (ANR), under grant ANR-23-CE31-0011 (project PEGaSUS). R.B., S.C., P.D.A., M.F., A.M., A.Ro., A.Re. acknowledge financial support from the GRAWITA Large Program Grant (PI P. D’Avanzo) and the PRIN-INAF 2022 ”Shedding light on the nature of gap transients: from the observations to the models”. R.B., S.C., P.D.A., M.F., A.M. acknowledge financial support from the Italian Space Agency, contract ASI/INAF No. I/004/11/6. A.Rossi acknowledges support from the INAF project Premiale Supporto Arizona & Italia. E.W.L. is supported by the National Natural Science Foundation of China (Grant No. 12133003). X.G.W. is supported by the National Natural Science Foundation of China (Grant No. 12373042) and the Bagui Scholars Program (No. GXR-6BG2424001). D.B.M. is funded by the European Union (ERC, HEAVYMETAL, 101071865, Views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them). The Cosmic Dawn Center (DAWN) is funded by the Danish National Research Foundation under grant No. DNRF140. Z.P.J. is supported by the National Natural Science Foundation of China (Grant Nos. 12225305 and 12321003). J.R. is supported by the General Fund (Grant No. 2024M763530) of the China Postdoctoral Science Foundation. We acknowledge the following observational support for this work: partly collected at the Schmidt Telescope of the INAF – Osservatorio Astronomico di Padova (Asiago, Italy) and with the Italian Telescopio Nazionale Galileo (TNG) operated on La Palma by the Fundación Galileo Galilei of the INAF at the Observatorio del Roque de los Muchachos of the Instituto de Astrofísica de Canarias; partly collected at the European Organisation for Astronomical Research in the Southern Hemisphere under ESO programme 110.24CF (PI Tanvir, Vergani, Malesani), and with the Nordic Optical Telescope (program 69-023; PI: Malesani), owned in collaboration by the University of Turku and Aarhus University, and operated jointly by Aarhus University, the University of Turku, and the University of Oslo (representing Denmark, Finland and Norway), the University of Iceland and Stockholm University; partly based on observations made at Observatoire de Haute Provence (CNRS), France, with MISTRAL. We also acknowledge the support from the staff at Nanshan Station of Xinjiang Astronomical Observatory (XAO) for their assistance with the Photometric Auxiliary Telescope (PAT-17). This research has made use of the MISTRAL database, operated at CeSAM (LAM), Marseille, France. This work made use of data supplied by the UK Swift Science Data Centre at the University of Leicester. We acknowledge the use of the Swift and Fermi archive’s public data.Appendix A
In this appendix, we provide detailed information on the observation data and fitting results of afterglow.
@@
78 | 9.8 | 13.61 | 0.07 | V |
105 | 19.7 | 15.64 | 0.04 | WHITE |
125 | 19.7 | 15.93 | 0.04 | WHITE |
145 | 19.7 | 16.26 | 0.04 | WHITE |
165 | 19.7 | 16.45 | 0.04 | WHITE |
185 | 19.7 | 16.59 | 0.05 | WHITE |
205 | 19.7 | 16.77 | 0.05 | WHITE |
225 | 19.7 | 16.97 | 0.05 | WHITE |
238 | 6.9 | 16.99 | 0.08 | WHITE |
346 | 78.7 | 17.72 | 0.08 | U |
426 | 78.7 | 18.01 | 0.09 | U |
510 | 87.6 | 17.95 | 0.08 | U |
570 | 18.7 | 17.45 | 0.18 | B |
594 | 18.7 | 18.38 | 0.10 | WHITE |
644 | 18.7 | 17.12 | 0.26 | V |
718 | 18.7 | 18.78 | 0.28 | U |
743 | 17.7 | 17.77 | 0.22 | B |
767 | 18.7 | 18.62 | 0.12 | WHITE |
817 | 17.7 | 17.06 | 0.26 | V |
1048 | 18.7 | 17.42 | 0.32 | V |
1121 | 18.7 | 18.74 | 0.31 | U |
1170 | 18.7 | 19.11 | 0.18 | WHITE |
@@
849 | 17.00 | 0.03 | clear | HMT | |
925 | 17.16 | 0.04 | clear | HMT | |
1001 | 17.26 | 0.04 | clear | HMT | |
1077 | 17.35 | 0.04 | clear | HMT | |
1154 | 17.37 | 0.05 | clear | HMT | |
1230 | 17.42 | 0.05 | clear | HMT | |
1306 | 17.48 | 0.05 | clear | HMT | |
1383 | 17.51 | 0.06 | clear | HMT | |
1459 | 17.60 | 0.06 | clear | HMT | |
1536 | 17.73 | 0.08 | clear | HMT | |
1613 | 17.71 | 0.08 | clear | HMT | |
1689 | 17.65 | 0.08 | clear | HMT | |
1766 | 17.80 | 0.08 | clear | HMT | |
1842 | 17.80 | 0.08 | clear | HMT | |
1919 | 17.97 | 0.09 | clear | HMT | |
1995 | 18.00 | 0.09 | clear | HMT | |
2071 | 17.91 | 0.08 | clear | HMT | |
2147 | 17.95 | 0.08 | clear | HMT | |
2224 | 18.00 | 0.09 | clear | HMT | |
2301 | 18.03 | 0.09 | clear | HMT | |
2377 | 17.94 | 0.08 | clear | HMT | |
2454 | 18.12 | 0.09 | clear | HMT | |
2530 | 18.33 | 0.12 | clear | HMT | |
2607 | 18.26 | 0.12 | clear | HMT | |
2683 | 18.28 | 0.14 | clear | HMT | |
2759 | 18.46 | 0.15 | clear | HMT | |
2836 | 18.40 | 0.16 | clear | HMT | |
2912 | 18.29 | 0.13 | clear | HMT | |
2989 | 18.30 | 0.12 | clear | HMT | |
3155 | 18.41 | 0.11 | clear | HMT | |
3292 | 18.33 | 0.10 | clear | HMT | |
3428 | 18.41 | 0.11 | clear | HMT | |
3565 | 18.34 | 0.10 | clear | HMT | |
3701 | 18.51 | 0.13 | clear | HMT | |
3838 | 18.49 | 0.14 | clear | HMT | |
3974 | 18.71 | 0.14 | clear | HMT | |
4111 | 18.59 | 0.11 | clear | HMT | |
4247 | 18.70 | 0.12 | clear | HMT | |
4384 | 18.63 | 0.11 | clear | HMT | |
4615 | 18.78 | 0.13 | clear | HMT | |
4832 | 18.88 | 0.14 | clear | HMT | |
5048 | 18.82 | 0.12 | clear | HMT | |
5264 | 18.86 | 0.11 | clear | HMT | |
5481 | 19.06 | 0.14 | clear | HMT | |
5776 | 19.16 | 0.14 | clear | HMT | |
6092 | 18.78 | 0.10 | clear | HMT | |
6409 | 18.90 | 0.11 | clear | HMT | |
6725 | 19.35 | 0.19 | clear | HMT | |
5647.5 | 18.88 | 0.03 | clear | T80 | |
6585.0 | 18.98 | 0.03 | clear | T80 | |
7523.5 | 19.21 | 0.04 | clear | T80 | |
8617.5 | 19.29 | 0.04 | clear | T80 | |
9868.0 | 19.49 | 0.05 | clear | T80 | |
12253.0 | 19.87 | 0.08 | clear | T80 | |
13889.5 | 20.02 | 0.08 | clear | T80 | |
15507.0 | 19.76 | 0.09 | clear | T80 | |
17228.0 | 19.82 | 0.13 | clear | T80 | |
19115.5 | 20.06 | 0.16 | clear | T80 | |
21063.5 | 20.28 | 0.24 | clear | T80 |
@@
37839.7 | 21.55 | 0.05 | g | VLT/XSHOOTER | |
37605.6 | 20.78 | 0.04 | r | VLT/XSHOOTER | |
38075.6 | 19.77 | 0.04 | z | VLT/XSHOOTER | |
40787.7 | 21.47 | 0.07 | g | NOT | |
41488.4 | 20.90 | 0.03 | r | NOT | |
122342.8 | 21.82 | 0.12 | r | NOT | |
41175.4 | 20.36 | 0.03 | i | NOT | |
41873.7 | 19.92 | 0.04 | z | NOT | |
123025.7 | 20.99 | 0.11 | z | NOT | |
205510.8 | 21.41 | 0.13 | z | NOT | |
558187.0 | 21.63 | 0.10 | z | NOT | |
36686. | 30. | 20.75 | 0.03 | r | GTC/OSIRIS+ |
36803. | 30. | 20.70 | 0.06 | r | GTC/OSIRIS+ |
36943. | 30. | 20.76 | 0.04 | r | GTC/OSIRIS+ |
900109. | 4250 | 22.99 | 0.04 | r | GTC/OSIRIS+ |
900873. | 1590 | 21.82 | 0.09 | z | GTC/OSIRIS+ |
976725. | 1590 | 21.45 | 0.03 | z | GTC/OSIRIS+ |
978336. | 4250 | 22.51 | 0.02 | r | GTC/OSIRIS+ |

### v2

@@ 3.2 Parameter inference
The multiband observational data and the model predictions based on the best-fit parameters are shown in Figure 7. The FS+RS afterglow model provides a good explanation for the optical data and is roughly consistent with the GeV observations (despite large uncertainties), but it does not explain the X-ray observations well, suggesting that there may be an additional contribution to the X-ray emission. By combining the afterglow model and prompt emission to constrain the initial Lorentz factor, we can examine the tight parameter correlation proposed by Liang et al. (2015), i.e. . As illustrated in the left side of Figure 8, GRB 240825A generally aligns with this correlation and, similar to other GRBs with RS emission (hereafter referred to as RS GRBs) (Japelj et al. 2014), exhibits higher luminosity compared to other typical Type II GRBs. Furthermore, based on the microphysics parameters of magnetic energy in the FS and RS obtained from the posterior parameters, we can estimate the magnetic ratio of RS and FS jet with 171717 is defined in Zhang & Kobayashi (2005). To compare with some statistical studies (Japelj et al. 2014; Yi et al. 2020; Wang et al. 2024), we adopt here, which requires taking the square root for conversion to the magnetic field ratio.. As shown on the right side of Figure 8, GRB 240825A exhibits a RS-to-FS magnetic ratio that is in the high end of the distribution derived by (Wang et al. 2024), and the of the host galaxy we finally obtained as free parameters matches well the constraint obtained by Cheng et al. (2025).
@@ 4 Summary and Discussion
In this work, we performed a detailed analysis of the prompt emission spectrum using data from Swift-BAT and Fermi-GBM/LAT and modeled the multiwavelength (optical/NIR, X-ray, -ray) afterglow light curves using the FS+RS model. By combining the observational data with our analysis results, we summarize the characteristics of this event as follows:
The prompt emission spectrum of GRB 240825A is best fit by a model combining a Band function and a power-law component, with a high-energy cutoff of 76 MeV. The peak energy () of the Band function is approximately 500 keV, and the -corrected isotropic energy is estimated to be 1.7 erg. These values make GRB 240825A consistent with the Amati relation.
Considering the high-energy cutoff of the prompt emission due to absorption and the radiation efficiency ( 0.1), constraints yield an initial Lorentz factor 245 and an isotropic kinetic energy 1.7 erg. These two constraints from prompt emission were used to update the prior ranges of parameters during Bayesian inference in the afterglow modeling.
@@
Based on the aforementioned characteristics of GRB 240825A, we have conducted discussions on the following points. Firstly, GRB 240825A is an event with comprehensive observational coverage across time and wavelength, spanning from the prompt emission to the afterglow phase. During the afterglow phase, the emission is initially dominated by RS components, and transitions to being FS dominated, and eventually reflects contributions from the host galaxy. Secondly, such a “flattening case” in the combination of FS+RS typically requires 1 (Zhang et al. 2003), which is consistent with the result obtained from our model fitting. This suggests that the RS is more magnetized than the FS, but is not Poynting-flux-dominated (otherwise the RS is suppressed). Secondly, although this burst appears to be a typical Type II GRB based on empirical correlations, its comprehensive observational data make it an excellent probe for studying the physics of GRB jets. For example, the initial Lorentz factor of the afterglow is often correlated with the onset time. Although SVOM/C-GFT conducted rapid follow-up observations of GRB 240825A, the optical rise phase is still missing. Therefore, the constraints on the Lorentz factor derived from the prompt emission are crucial, highlighting the importance of comprehensive observations from the prompt emission to the afterglow phase. Thirdly, despite the overall consistency between the model and the observed multiwavelength afterglow data, the X-ray light curve is not well reproduced in our current framework. This discrepancy may point to additional physical components or emission mechanisms not accounted for in the standard external shock model. Several scenarios have been proposed in the literature to explain the shaping of X-ray afterglow emission in GRBs, such as energy injection, evolving shock parameters, etc. (Zhang et al. 2006; Fan & Piran 2006). However, most of these scenarios generally predict correlated features in other bands, particularly in the optical or high-energy regimes. The fact that our model provides a good fit to the optical, while significantly underpredicting the X-ray flux, suggests that the X-ray excess may arise from an additional component that selectively contributes to the X-ray band. In previous work, such as GRB 130427A, X-rays may come from the contributions of flash-RS and RS Vestrand et al. (2014). In addition, based on GRB 221009A, people realized that there may be an extremely narrow component in the jet structure Zhang et al. (2024); Ren et al. (2024), i.e., the two-component jet model. Hence, this remains an open question, and future observations and modeling efforts will be necessary to fully understand the origin of such X-ray behavior.
@@ Acknowledgements.
The Space-based multiband Variable Objects Monitor (SVOM) is a joint Chinese-French mission led by the Chinese National Space Administration (CNSA), the French Space Agency (CNES), and the Chinese Academy of Sciences (CAS). We gratefully acknowledge the unwavering support of NSSC, IAMCAS, XIOPM, NAOC, IHEP, CNES, CEA, and CNRS. This work is supported by the National Key R&D Program of China (grant No. 2024YFA1611600), the SVOM project (a mission under the Strategic Priority Program on Space Science of the Chinese Academy of Sciences), the Strategic Priority Research Program of the Chinese Academy of Sciences (NFSC, grant No. XDB0550401), and the National Natural Science Foundation of China (NSFC, grant No. 12494573). This research was partly supported by Natural Science Foundation of Xinjiang Uygur Autonomous Region (grant No. 2024D01D32), Tianshan Talent Training Program (grant No. 2023TSYCLJ0053), and Tianshan Innovation Team Program (grant No. 2024D14015). This work is also supported by the following funding programs: Y.W. is supported by the Jiangsu Funding Program for Excellent Postdoctoral Talent (grant No. 2024ZB110), the Postdoctoral Fellowship Program (grant No. GZC20241916), and the General Fund (grant No. 2024M763531) of the China Postdoctoral Science Foundation. G.P.L. is supported by a Royal Society Dorothy Hodgkin Fellowship (grant Nos. DHF-R1-221175 and DHF-ERE-221005). A.S. acknowledges support by a postdoctoral fellowship from the CNES. X.H.H. is supported by the National Key R&D Program of China (grant No. 2024YFA1611702) and the Strategic Priority Research Program of the Chinese Academy of Sciences (grant No. XDB0550101). S.D.V., E.L.F., B.S. acknowledge the support of the French Agence Nationale de la Recherche (ANR), under grant ANR-23-CE31-0011 (project PEGaSUS). R.B., S.C., P.D.A., M.F., A.M., A.Ro., A.Re. acknowledge financial support from the GRAWITA Large Program grant (PI P. D’Avanzo) and the PRIN-INAF 2022 ”Shedding light on the nature of gap transients: from the observations to the models”. R.B., S.C., P.D.A., M.F., A.M. acknowledge financial support from the Italian Space Agency, contract ASI/INAF No. I/004/11/6. A.Rossi acknowledges support from the INAF project Premiale Supporto Arizona & Italia. E.W.L. is supported by the National Natural Science Foundation of China (NSFC, grant No. 12133003). X.G.W. is supported by the National Natural Science Foundation of China (NSFC, grant No. 12373042) and the Bagui Scholars Program (No. GXR-6BG2424001). D.B.M. is funded by the European Union (ERC, HEAVYMETAL, 101071865, Views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them). The Cosmic Dawn Center (DAWN) is funded by the Danish National Research Foundation under grant No. DNRF140. Z.P.J. is supported by the National Natural Science Foundation of China (NSFC, grant Nos. 12225305 and 12321003). D.M.W.is supported by the National Natural Science Foundation of China (NSFC, grant No. 12473049. J.R. is supported by the General Fund (grant No. 2024M763530) of the China Postdoctoral Science Foundation. We acknowledge the following observational support for this work: partly collected at the Schmidt Telescope of the INAF – Osservatorio Astronomico di Padova (Asiago, Italy) and with the Italian Telescopio Nazionale Galileo (TNG) operated on La Palma by the Fundación Galileo Galilei of the INAF at the Observatorio del Roque de los Muchachos of the Instituto de Astrofísica de Canarias; partly based on data obtained with the instrument OSIRIS, built by a Consortium led by the Instituto de Astrofísica de Canarias in collaboration with the Instituto de Astronomía of the Universidad Autónoma de México. OSIRIS was funded by GRANTECAN and the National Plan of Astronomy and Astrophysics of the Spanish Government (program: GTCMULTIPLE2F-24A, PI: Agüí Fernández); partly collected at the European Organisation for Astronomical Research in the Southern Hemisphere under ESO programme 110.24CF (PI Tanvir, Vergani, Malesani), and with the Nordic Optical Telescope (program 69-023; PI: Malesani), owned in collaboration by the University of Turku and Aarhus University, and operated jointly by Aarhus University, the University of Turku, and the University of Oslo (representing Denmark, Finland and Norway), the University of Iceland and Stockholm University; partly based on observations made at Observatoire de Haute Provence (CNRS), France, with MISTRAL. We also acknowledge the support from the staff at Nanshan Station of Xinjiang Astronomical Observatory (XAO) for their assistance with the Photometric Auxiliary Telescope (PAT-17). This research has made use of the MISTRAL database, operated at CeSAM (LAM), Marseille, France. This work made use of data supplied by the UK Swift Science Data Centre at the University of Leicester. We acknowledge the use of the Swift and Fermi archive’s public data.Appendix A Afterglow observation data and spectral fitting results
In this appendix, we provide detailed information on the observation data and fitting results of afterglow. The observation data from various telescopes are listed as follows: SVOM C-GFT (Table LABEL:tab_cgft), SVOM VT (Table 3), Swift UVOT (Table 4), HMT/T80 (Table LABEL:tab_xm), PAT17/ALT100C (Table LABEL:tab_xm2), LCOGT (Table 7), VLT/NOT/GTC (Table 8), and Asiago/REM/TNG/LBT/MISTRAL (Table 9). Table 10 presents the priors and posteriors of the afterglow model parameters. Figure 9 shows the corner plot of the posterior distributions for the afterglow fitting, and the corresponding parameter values, along with their prior ranges, are listed in Table 10.
@@
| 78 | 9.8 | 13.61 | 0.07 | V |
| 105 | 19.7 | 15.64 | 0.04 | WHITE |
| 125 | 19.7 | 15.93 | 0.04 | WHITE |
| 145 | 19.7 | 16.26 | 0.04 | WHITE |
| 165 | 19.7 | 16.45 | 0.04 | WHITE |
| 185 | 19.7 | 16.59 | 0.05 | WHITE |
| 205 | 19.7 | 16.77 | 0.05 | WHITE |
| 225 | 19.7 | 16.97 | 0.05 | WHITE |
| 238 | 6.9 | 16.99 | 0.08 | WHITE |
| 346 | 78.7 | 17.72 | 0.08 | U |
| 426 | 78.7 | 18.01 | 0.09 | U |
| 510 | 87.6 | 17.95 | 0.08 | U |
| 570 | 18.7 | 17.45 | 0.18 | B |
| 594 | 18.7 | 18.38 | 0.10 | WHITE |
| 644 | 18.7 | 17.12 | 0.26 | V |
| 718 | 18.7 | 18.78 | 0.28 | U |
| 743 | 17.7 | 17.77 | 0.22 | B |
| 767 | 18.7 | 18.62 | 0.12 | WHITE |
| 817 | 17.7 | 17.06 | 0.26 | V |
| 1048 | 18.7 | 17.42 | 0.32 | V |
| 1121 | 18.7 | 18.74 | 0.31 | U |
| 1170 | 18.7 | 19.11 | 0.18 | WHITE |
@@
| 849 | 17.00 | 0.03 | clear | HMT | |
| 925 | 17.16 | 0.04 | clear | HMT | |
| 1001 | 17.26 | 0.04 | clear | HMT | |
| 1077 | 17.35 | 0.04 | clear | HMT | |
| 1154 | 17.37 | 0.05 | clear | HMT | |
| 1230 | 17.42 | 0.05 | clear | HMT | |
| 1306 | 17.48 | 0.05 | clear | HMT | |
| 1383 | 17.51 | 0.06 | clear | HMT | |
| 1459 | 17.60 | 0.06 | clear | HMT | |
| 1536 | 17.73 | 0.08 | clear | HMT | |
| 1613 | 17.71 | 0.08 | clear | HMT | |
| 1689 | 17.65 | 0.08 | clear | HMT | |
| 1766 | 17.80 | 0.08 | clear | HMT | |
| 1842 | 17.80 | 0.08 | clear | HMT | |
| 1919 | 17.97 | 0.09 | clear | HMT | |
| 1995 | 18.00 | 0.09 | clear | HMT | |
| 2071 | 17.91 | 0.08 | clear | HMT | |
| 2147 | 17.95 | 0.08 | clear | HMT | |
| 2224 | 18.00 | 0.09 | clear | HMT | |
| 2301 | 18.03 | 0.09 | clear | HMT | |
| 2377 | 17.94 | 0.08 | clear | HMT | |
| 2454 | 18.12 | 0.09 | clear | HMT | |
| 2530 | 18.33 | 0.12 | clear | HMT | |
| 2607 | 18.26 | 0.12 | clear | HMT | |
| 2683 | 18.28 | 0.14 | clear | HMT | |
| 2759 | 18.46 | 0.15 | clear | HMT | |
| 2836 | 18.40 | 0.16 | clear | HMT | |
| 2912 | 18.29 | 0.13 | clear | HMT | |
| 2989 | 18.30 | 0.12 | clear | HMT | |
| 3155 | 18.41 | 0.11 | clear | HMT | |
| 3292 | 18.33 | 0.10 | clear | HMT | |
| 3428 | 18.41 | 0.11 | clear | HMT | |
| 3565 | 18.34 | 0.10 | clear | HMT | |
| 3701 | 18.51 | 0.13 | clear | HMT | |
| 3838 | 18.49 | 0.14 | clear | HMT | |
| 3974 | 18.71 | 0.14 | clear | HMT | |
| 4111 | 18.59 | 0.11 | clear | HMT | |
| 4247 | 18.70 | 0.12 | clear | HMT | |
| 4384 | 18.63 | 0.11 | clear | HMT | |
| 4615 | 18.78 | 0.13 | clear | HMT | |
| 4832 | 18.88 | 0.14 | clear | HMT | |
| 5048 | 18.82 | 0.12 | clear | HMT | |
| 5264 | 18.86 | 0.11 | clear | HMT | |
| 5481 | 19.06 | 0.14 | clear | HMT | |
| 5776 | 19.16 | 0.14 | clear | HMT | |
| 6092 | 18.78 | 0.10 | clear | HMT | |
| 6409 | 18.90 | 0.11 | clear | HMT | |
| 6725 | 19.35 | 0.19 | clear | HMT | |
| 5647.5 | 18.88 | 0.03 | clear | T80 | |
| 6585.0 | 18.98 | 0.03 | clear | T80 | |
| 7523.5 | 19.21 | 0.04 | clear | T80 | |
| 8617.5 | 19.29 | 0.04 | clear | T80 | |
| 9868.0 | 19.49 | 0.05 | clear | T80 | |
| 12,253.0 | 19.87 | 0.08 | clear | T80 | |
| 13,889.5 | 20.02 | 0.08 | clear | T80 | |
| 15,507.0 | 19.76 | 0.09 | clear | T80 | |
| 17,228.0 | 19.82 | 0.13 | clear | T80 | |
| 19,115.5 | 20.06 | 0.16 | clear | T80 | |
| 21,063.5 | 20.28 | 0.24 | clear | T80 |
@@
| 37,839.7 | 21.55 | 0.05 | g | VLT/XSHOOTER | |
| 37,605.6 | 20.78 | 0.04 | r | VLT/XSHOOTER | |
| 38,075.6 | 19.77 | 0.04 | z | VLT/XSHOOTER | |
| 40,787.7 | 21.47 | 0.07 | g | NOT | |
| 41,488.4 | 20.90 | 0.03 | r | NOT | |
| 122,342.8 | 21.82 | 0.12 | r | NOT | |
| 41,175.4 | 20.36 | 0.03 | i | NOT | |
| 41,873.7 | 19.92 | 0.04 | z | NOT | |
| 123,025.7 | 20.99 | 0.11 | z | NOT | |
| 205,510.8 | 21.41 | 0.13 | z | NOT | |
| 558,187.0 | 21.63 | 0.10 | z | NOT | |
| 36,686. | 30. | 20.75 | 0.03 | r | GTC/OSIRIS+ |
| 36,803. | 30. | 20.70 | 0.06 | r | GTC/OSIRIS+ |
| 36,943. | 30. | 20.76 | 0.04 | r | GTC/OSIRIS+ |
| 900,109. | 4250 | 22.99 | 0.04 | r | GTC/OSIRIS+ |
| 900,873. | 1590 | 21.82 | 0.09 | z | GTC/OSIRIS+ |
| 976,725. | 1590 | 21.45 | 0.03 | z | GTC/OSIRIS+ |
| 978,336. | 4250 | 22.51 | 0.02 | r | GTC/OSIRIS+ |